OpenStudy (anonymous):

The population (in millions) of a certain island can be approximated by the function p(x)=50(1.05)^x, where x is the number of years since 2000. In which year will the population reach 200 million? Hint: An answer such as 2002.4 would represent the year 2002. A. 2028 B. 2066 C. 2002 D. 2015

OpenStudy (anonymous):

@kaylamarie101

OpenStudy (anonymous):

200,000,000=50(1.05)^x then i divided each side by 50 which gave me 4,000,000= 1.05^x then log_1.05 4,000,000 = x x=311

OpenStudy (anonymous):

after that im not sure what to do @kaylamarie101

OpenStudy (anonymous):

Okay. give me a moment

OpenStudy (anonymous):

sure :) Im not sure if my work is correct

OpenStudy (anonymous):

Okay. Since you said x=311, you use the function and substitute every x as 311

OpenStudy (anonymous):

Okay, first, you did it wrong..You but the 200 million as p(x) which is not correct

OpenStudy (anonymous):

ahhh darn it :O

OpenStudy (anonymous):

The purpose of this is to find out in which year the population will be 200 million. I see where you might have thought that that would be right, but let me explain it to you.

OpenStudy (anonymous):

As an example, lets just pick on of the choices. Im going to pick 2028 because it sounds cool.

OpenStudy (anonymous):

It says "p(x)=50(1.05)^x, where x is the number of years since 2000."

OpenStudy (anonymous):

So, the key is seeing how many years it has been since 2000 because that will be x. So, Tell me, in 2028, how many years has it been since 2000?

OpenStudy (anonymous):

@joanne2014 ? Can you tell me that?

OpenStudy (anonymous):

yea im figuring it out i think

OpenStudy (anonymous):

It is just subtraction hun. 2028-2000 will give you the difference between the two years.

OpenStudy (anonymous):

haha i figured at the end lol its 28

OpenStudy (anonymous):

exactly. Which means that x=28. Correct?

OpenStudy (anonymous):

yea so then we plug 28 in for the exponent or p(x)?

OpenStudy (anonymous):

Both. whether f(x), g(x), h(x), or p(x) when you find out what x is, you substitute the x for the number. So, in this case, x=28 and we have p(x). So: p(28)=5(1.05)^28 Tell me what that is?

OpenStudy (anonymous):

is it 19.6?

OpenStudy (anonymous):

Correct! Now, is that 200 million?

OpenStudy (anonymous):

no so i keep looking at the years until it equals 200 million right?

OpenStudy (anonymous):

Exactly ! (: You Continue to find the difference of the year you chose and 2000 and plug that number into x :D

OpenStudy (anonymous):

You are doing really good! :)

OpenStudy (anonymous):

yay thankk you !

OpenStudy (anonymous):

No problem Darlin'! I'm here when you need me!

OpenStudy (anonymous):

And thanks for the medal!

OpenStudy (anonymous):

Thanks for being patient :D